OpenStudy (anonymous):

Find the local maximum and minimum values of f using both the First and Second Derivative Tests. f(x)= x/(x^(2)+16). I got y=4 for local max and y=-4 for local minimum, but it says it's wrong. Not sure what I'm doing wrong.

OpenStudy (anonymous):

first, write your f'(x) hint: you'd have to se the quotient rule

OpenStudy (anonymous):

f'(x)= (x^(2)+16)*1-x*2x/(x^(2)+16)^(2) => x^(2)+16-2x^(2)/(x^(2)+16)^(2) =>16-x^(2)/(x^(2)+16)^(2) and I set x^(2)/(x^(2)+16)^(2) =0.

OpenStudy (anonymous):

you can use the equation editor button below so people can understand what you just wrote :)

OpenStudy (anonymous):

\[f'(x)= (x^2+16)*1-x*2x/(x^2+16)^2\]

OpenStudy (anonymous):

\[x^2+16-2x^2/(x^2+16)^2 \]

OpenStudy (anonymous):

\[f'(x): 16-x^2/(x^2+16)^2=0\]

OpenStudy (anonymous):

so, \[16-x^2=0\implies x=\pm4 \]

OpenStudy (anonymous):

one point should be the maxima and the other, the minima we check using second derivative. similarly, express your second derivative f''(x)=?

OpenStudy (anonymous):

\[-2x^3-32x-64x+4x^3/(x^2+9)^3\]

OpenStudy (anonymous):

Refer to the attachment. Calculations were performed by Mathematica.